No hand writing please

On September 30, 2015, Ericson Company negotiated a two-year, 1,200,000 dudek loan from a foreign bank at an interest rate ofa. Prepare all journal entries related to this foreign currency borrowing assuming the above exchange rates for 1 dudek. (DoView transaction list X: 1 Record the note and conversion of 1,200,000 million dudeks into $ at the spot rate. Record the accView transaction list record the foreign exchange gain/loss thereof. 4 Record the first annual interest payment including anyb. Determine the effective cost of borrowing in dollars in each of the three years 2015, 2016, and 2017. (Do not round interm
Answer #1

a)

Date General Journal Debit Credit
09/30/2015 Cash 144,000
Note Payable 144,000
(1,200,000*0.120)
12/31/2015 Interest Expense 750
Interest Payable 750
[1,200,000*2%*(3/12) ]=6,000dudeks *0.125
12/31/2015 Foreign Exchange Loss 6,000
Note payable 6,000
[1,200,000 * (0.125-0.120)]
9/30/2016 Interest expense 2,520
Interest payable 750
Foreign exchange loss 90
Cash 3,360
Foreign exchange loss = 6,000*(0.140-0.125) =
Interest Expense = [1,200,000*2%*(9/12)]=18,000dudeks*0.140 =
12/31/2016 Interest Expense 870
Interest Payable 870
(6,000 dudeks * 0.145)
12/31/2016 Foreign exchange loss 24,000
Note Payable 24,000
[1,200,000 * (0.145-0.125)]
9/30/2017 Interest Expense 3,060
Interest Payable 870
Foreign exchange loss 150
Cash 4,080
Interest expense [18,000dudeks *0.170] = 3,060
Foreign exchange loss[6,000 dudeks * (0.170-0.145)]
Cash [24,000 dudeks * 0.170] =
9/30/2017 Note Payable 174,000
Foreign exchange loss 30,000
Cash 204,000
Foreign exhange loss = 6,000 + 24,000 = 30,000
Cash = [1,200,000 * 0.170] = 204,000

b) Effective cost of borrowing:

2015

Interest Expense = 750

Foreign exhange loss = 6,000

TOTAL = 6,750

(6,750 / 144,000)*(12/3) *100 = 18.75% for 12 months

2016

Interest Expense = 3,390 [2,520 + 870]

Foreign exchange loss = 24,090 [24,000 + 90]

TOTAL = 27,480

(27,480 / 144,000) * 100 = 19.08% for 12 months

2017:

Interest Expense = 3,060

Foreign exchange loss = 30,150 [30,000 + 150]

TOTAL = 33,210

(33,210 / 144,000) * (12/9) * 100 = 30.75% for 12 months